# The Impact of Responsible AI Research on Innovation and Development

### v1

@@ Introduction
We acknowledge six limitations. First, the total number of papers analyzed may not represent the full extent of published papers as we relied on Semantic Scholar. Nevertheless, Semantic Scholar has been used in similar analyses (Cao et al. 2023) with coverage comparable to Google Scholar (Hannousse 2021). To mitigate potential biases, future research could replicate our method using alternative databases such as Crossref (cro 1999) and OpenAlex (ope 2023). Second, our title-matching method for associating Semantic Scholar papers with patent references may not be exhaustive. Future studies could explore alternative matching algorithms such as Biblio Glutton (bib 2022) and Marx and Fuegi (2022)’s. Third, our approach does not capture the intent for a citation. When we attempted to identify specific intents by searching for mentions of particular unigrams such as “extension”, “future”, and “use” in citations (similar to what was done in previous work (Cohan et al. 2019)), we encountered two main challenges: 1) establishing a ground truth was difficult and relied on our estimations of potential intent; and 2) even with a preliminary ground truth, the results were not promising (F1 = 0.67 for intent classification as reported in (Cohan et al. 2019)). To this end, we deliberately decided not to capture the intent for citations. Fourth, as we shall see next, our metrics for assessing research impact are limited to data from USPTO (patents) and GitHub (repositories). Future studies could expand to a wider range of sources (e.g., patents from the European Patent Office), enhancing the comprehensiveness of impact assessment. Fifth, as RAI is an emerging area, we acknowledge that the number of papers may currently be limited; however, this can change drastically in the future. Finally, as we shall see next, our metrics have limitations, particularly in capturing broader societal impact. Future work could investigate how research translates into legislation by comparing RAI paper topics with legislative drafting such as the EU AI Act, or by examining whether research adequately covers real-world incidents by comparing RAI paper topics with databases collating AI incidents (AIID 2024).
@@
Table 1: Analysis variables. We used these variables and notations throughout our analysis to measure the impact of Responsible AI research on innovation (through patents) and development (through code repositories).
@@
$P$ Set of patents in the corpus (1980-2022, USPTO).
$R$ Set of repositories in the corpus (2008-2022, GitHub).
@@
$U$ Set of papers in the corpus (2015-2022, written in the English language).
$H$ Set of papers from the venues under study.
$H^{\prime}$ Set of papers not from the venues under study.
$c_{\mathrm{in}}(y)$ Set of papers associated with a document $y$ (another paper or a patent).
$c_{\mathrm{out}}(u)$ Set of documents (among papers and patents) citing paper $u$.
$H\mid X$ Set of papers in $H$ that went into patents or into repositories.
$(H\mid X)^{\prime}$ Set of papers in $H$ that did not go into patents or go into repositories.
$H^{\prime}\mid X$ Set of papers in $H^{\prime}$ that went into patents or into repositories.
$X\mid H$ Set of patents or repositories that cite papers in $H$.
@@
Author Positionality and Ethical Considerations
The research team includes three men from Southeast Asia and Southern Europe, working in an industrial research lab. Our shared backgrounds include Human-Computer Interaction, software engineering, AI, social and ubiquitous computing, and urbanism.
In light of recent calls for Responsible AI research to include impact and ethical statements (Olteanu et al. 2023), our research neither involves any subjects nor do we foresee any significant harm that could result.
@@ Results
Figure 3: Number of papers for each RAI topic over the analyzed period (2015-2022). Since 2017, there has been an increasing trend of RAI papers, especially in Fairness and Privacy.
@@
Two potential reasons might explain this lag. The first is related to the longer process of publishing patents compared to papers. This leads to a situation where it might take up to a decade for a paper to accumulate patent citations (Cao et al. 2023). The frequency with which a paper is cited in a patent is influenced by the time elapsed since its publication. To address this time dependency, we conducted a survival analysis, as detailed in §Metrics. As observed in Figure 4a, the probability of a paper going into patents increases with each passing year after its publication. However, there could be another reason. It might also be due to the overproduction of papers on certain topics (e.g., Fairness or Privacy), which exceeds the rate of patent production. In simpler terms, if a topic becomes overly popular, it might lead to an oversaturation of papers, thus lowering the probability of a paper going into patents. This can create a bottleneck effect, straining the already lengthy patenting process, which typically takes 24-30 months from filing to grant (usp 2015).
@@
Top institutions with papers going into patents and repositories are Cornell University and National University of Singapore (NUS), respectively. Carnegie Mellon University (CMU) produces the most RAI papers across all topics (Table 3). However, NUS is particularly prominent when considering papers going into repositories, with significant contributions in Fairness and Privacy. Overall, universities tend to have a greater impact on responsible AI than companies. However, when grouping the number of RAI papers by topic, we observed that companies were also leading in certain areas. In Accountability, Amazon and Google were among the top institutions, along with Stanford University and Oxford University; in Explainability, Microsoft shares the top places with TU Berlin; and in Fairness, Google leads along with NUS and CMU.
@@
RAI papers going into patents or repositories have unconventional combinations of prior work. In general, papers going into patents or repositories are grounded on unconventional combinations of prior work. However, Privacy papers that went into patents can be more conventional (Figure 5a), similar to Sustainability papers that went into repositories (Figure 5b). To see this distinction in a concrete example, consider the two most unconventional papers: “Alleviating Privacy Attacks via Causal Learning” by Tople, Sharma, and Nori (2020) and “Practical One-Shot Federated Learning for Cross-Silo Setting” by Li, He, and Song (2021). These papers draw from uncommon pairs of venues such as NeurIPS and IEEE Computer Security Foundations Symposium, and Conference on Computer and Communications Security and ICLR and have consequently gone into several repositories. On the other hand, we found that Sustainability papers going into repositories are more likely to be grounded on conventional combinations of prior work.
@@ Discussion
By analyzing 1,747 papers in the field of Responsible AI from leading Computer Science conferences, we found that a small subset that goes into patents or repositories is highly cited. The majority of these papers that go into patents focus on sustainability and privacy, while those that go into repositories focus on explainability. The temporal dynamics between papers that go into patents and those that go into repositories vary. The time frame for papers that go into patents is 8 years, while those that go into repositories are 1 year. Our survival analysis revealed not only the presence of a rapid publication cycle and a slower patent production cycle, but also that these cycles vary across different RAI topics. This suggests the need to reevaluate existing assumptions about the pace and consistency of knowledge production in both patents and repositories. When examining the top institutions producing impactful RAI papers in terms of patents and repositories, we observed that the impact of top U.S. institutions is not exclusive. Although our study concentrated on patents in the U.S., we noticed a significant contribution from European and Asian institutions. This indicates a more global distribution of influence than previously recognized. Finally, RAI papers generally rely on unconventional knowledge, often integrating expertise from fields such as Machine Learning (ML) and Human-Computer Interaction (HCI), in both patents and repositories.
@@
By synthesizing our findings with those documented in prior literature, we propose three recommendations for enhancing the translational landscape of Responsible AI research.
@@
Developing Joint Initiatives Between Academia and Industry. Translational research is crucial in bridging the gap between scientific breakthroughs and their practical application (Searles et al. 2016). Echoing the findings of previous studies (Cao et al. 2023), our study suggests that only a portion of Responsible AI (RAI) research significantly contributes to patents and repositories. Therefore, it is beneficial to foster collaborative efforts between academia and industry. This can be achieved through various initiatives. First, by establishing dual mentoring programs that connect RAI researchers with mentors from both academic and industrial sectors, including patent professionals. Second, by organizing joint conferences and patent clinics that unite AI researchers, patent engineers, and industry experts. Hosting patent clinics at conferences such as FAccT offers researchers immediate access to professional guidance on patenting, fostering a deeper exchange of ideas and encouraging patenting. Finally, by promoting corporate participation in academic endeavors through hackathons and sponsorships (Briscoe and Mulligan 2014), with the goal of enhancing the translation of RAI research into practical innovations.
@@
Balancing and Diversifying Focus in Responsible AI Research. The observed delay between RAI papers and patents may suggest an overemphasis on certain research topics. While this focus on specific areas (e.g., Fairness and Accountability) advances theoretical knowledge, it might inadvertently slow down the development of practical innovations (McKinsey & Company 2022). However, slowing down is not necessarily bad. In fact, incremental approaches could be favorable when considering the impact of technical work on social change (Abebe et al. 2020). This is advised because otherwise researchers may fail to recognize how changes in behaviors of pre-existing social systems can be caused by the introduction of technology. It can also lead to technical solutions focused purely ethical or value-aligned deployments, without considering whether a given system functions and provides any benefits at all (Raji et al. 2022), or even whether the best solution to a problem is a technical one (Selbst et al. 2019). To address these challenges, a more strategic approach is necessary to accelerate the transition from academic discoveries to patentable technologies. This would involve reevaluating the research priorities within the RAI community to ensure they align with practical needs, promoting a balance between academic exploration and outcomes that can be patented. Our findings also suggest that papers that go into patents tend to concentrate on sustainability and privacy, whereas those in repositories are focused on explainability. This discrepancy highlights the importance for academia to realign its research priorities with industry demands. One approach could be adopting a dynamic method for selecting topics, which remains responsive to the changing needs of the industry.
@@
Fostering Cross-discipline Responsible AI Research. The impact of RAI papers is shaped by how they build upon prior knowledge. Generally, RAI papers tend to use unconventional combinations of prior knowledge, and often draw upon expertise from diverse fields such as ML and HCI. To continue making a significant impact on both innovation and development, it is important to encourage researchers to engage in cross-disciplinary RAI research. This can be achieved by the formation of research teams that include experts from multiple disciplines (Rakova et al. 2021) (e.g., computer science, social sciences, ethics, and law), incorporating diverse perspectives and technical expertise (Moitra et al. 2022) as well as contributing towards organizational culture shift (Rakova et al. 2021). Additionally, training and educational programs that promote interdisciplinary work (Frodeman 2011) are required to train AI researchers in ethical, legal, and social implications of their work (McDonald and Pan 2020).
@@ Conclusion
By analyzing 1,747 RAI papers, we found that Responsible AI research, despite being an emerging field, contributes to patents and development. This translation process typically occurs within a time frame of 1 to 8 years, highlighting a gap between research output and translational outcomes. Our findings also challenge the conventional notion that impactful RAI research is predominantly the domain of top U.S. institutions. Instead, European and Asian institutions are also major contributors, indicating a more globally distributed expertise in the field. Finally, the multidisciplinary nature of RAI papers, which often build upon unconventional combinations of prior knowledge, is vital for ensuring the development of fair, transparent, and accountable AI systems.
@@ Appendix A Keywords to Filter RAI Papers
Removed keywords. These keywords were removed from our query because they were non-discriminative as they are often used to describe generally AI. The 19 removed keywords included: validation, reliability, correctness, accuracy, robustness, generalizability, authenticity, quality, measurability, dependability, capability, safety, security, resilience, confidentiality, integrity, availability, usability, controllability.
@@ Appendix B Analyzed Venues
Figure B.1: The cumulative number of citations in millions as a function of top venues considered in each research area. The top 10 venues in the six research areas (AI, Database, NLP, HCI, Data Mining, and Computer Vision) cover more than 88% of the total citations in each area. The numbers at the top of each subplot show the total coverage of citations from the top 10 venues in each area.
Table B.1: Distribution of papers (including non-RAI papers) over the years for each research area. Venues are ranked based on their h5 index impact factor. According to Google Scholar, h5 index is the h-index for articles published in the last five complete years; it is the largest number $h$ such that $h$ articles published in 2018-2022 have at least h citations each.

### v4

@@ Introduction
We acknowledge six limitations. First, the total number of papers analyzed may not represent the full extent of published papers as we relied on Semantic Scholar. Nevertheless, Semantic Scholar has been used in similar analyses (Cao et al. 2023) with coverage comparable to Google Scholar (Hannousse 2021). To mitigate potential biases, future research could replicate our method using alternative databases such as Crossref [^8] https://www.crossref.org and OpenAlex. [^9] https://openalex.org Second, our title-matching method for associating Semantic Scholar papers with patent references may not be exhaustive. Future studies could explore alternative matching algorithms such as Biblio Glutton [^10] https://github.com/kermitt2/biblio-glutton and Marx and Fuegi (2022)’s. Third, our approach does not capture the intent for a citation. When we attempted to identify specific intents by searching for mentions of particular unigrams such as “extension”, “future”, and “use” in citations (similar to what was done in previous work (Cohan et al. 2019)), we encountered two main challenges: 1) establishing a ground truth was difficult and relied on our estimations of potential intent; and 2) even with a preliminary ground truth, the results were not promising (F1 = 0.67 for intent classification as reported in (Cohan et al. 2019)). To this end, we deliberately decided not to capture the intent for citations. Fourth, as we shall see next, our metrics for assessing research impact are limited to data from USPTO (patents) and GitHub (repositories). Future studies could expand to a wider range of sources (e.g., patents from the European Patent Office), enhancing the comprehensiveness of impact assessment. Fifth, as RAI is an emerging area, we acknowledge that the number of papers may currently be limited; however, this can change drastically in the future. Finally, as we shall see next, our metrics have limitations, particularly in capturing broader societal impact. Future work could investigate the overlap between research studies that shape policy and those that influence patents. For example, one may examine how research translates into legislation by comparing RAI paper topics with legislative drafting such as the EU AI Act or the NIST’s AI RMF, or by examining whether research adequately covers real-world incidents by comparing RAI paper topics with databases collating AI incidents (AIID 2024). Future research could also replicate our methodology to study how authors’ diversity (e.g., nationalities, backgrounds, and areas of expertise) translates to varying levels of impact (Septiandri et al. 2023; Septiandri, Constantinides, and Quercia 2024).
@@
Table 1: We used these variables throughout our analysis to measure the impact of RAI research on innovation (through patents) and development (through code repositories).
@@
$P$ Patents in the corpus (1980-2022, USPTO).
$R$ Repositories in the corpus (2008-2022, GitHub).
@@
$U$ Papers in the corpus (2015-2022).
$H$ Papers from the venues under study.
$H^{\prime}$ Papers not from the venues under study.
$c_{\mathrm{in}}(y)$ Papers associated with a document $y$ (paper/patent).
$c_{\mathrm{out}}(u)$ Documents (among papers/patents) citing paper $u$.
$H\mid X$ Papers in $H$ that went into patents/repositories.
$(H\mid X)^{\prime}$ Papers in $H$ that didn’t go into patents/repositories.
$H^{\prime}\mid X$ Papers in $H^{\prime}$ that went into patents/repositories.
$X\mid H$ Patents or repositories that cite papers in $H$.
@@ Results
To provide context for these findings, we compared the percentage of RAI papers that went into patents or code repositories with similar metrics from other AI subfields such as Computer Vision and NLP. By computing our metrics on all papers in our dataset (whether RAI or not), we found that approximately 16.9% of papers in Computer Vision and 18.7% in NLP went into patents or code repositories and received 30.6% and 32.4% of academic citations, respectively. The lower rates for RAI papers compared to other AI subfields can be attributed either to the relatively young yet exponentially growing nature of RAI research, or to the possibility that more translational work is needed.
@@
Figure 3: Number of papers for each RAI topic between 2015-2022. Since 2017, there has been an increasing trend of RAI papers, especially in Fairness and Privacy.
@@
Two potential reasons might explain this lag. The first is related to the longer process of publishing patents compared to papers. This leads to a situation where it might take up to a decade for a paper to accumulate patent citations (Cao et al. 2023). The frequency with which a paper is cited in a patent is influenced by the time elapsed since its publication. To address time dependency, we conducted a survival analysis, as detailed in §Metrics. As observed in Figure 4a, the probability of a paper going into patents increases with each passing year after its publication. However, it might also be due to the overproduction of papers on certain topics (e.g., Fairness or Privacy), which exceeds the rate of patent production. In simpler terms, if a topic becomes overly popular, it might lead to an oversaturation of papers, thus lowering the probability of a paper going into patents. This can create a bottleneck effect, straining the already lengthy patenting process, which typically takes 24-30 months from filing to grant (Graham, Marco, and Miller 2015).
@@
Top institutions with papers going into patents and repositories are Cornell University and National University of Singapore (NUS), respectively. Carnegie Mellon University (CMU) produces the most RAI papers across all topics (Table 3). However, NUS is particularly prominent when considering papers going into repositories, with significant contributions in Fairness and Privacy. Overall, universities tend to have a greater impact on RAI than companies. However, when grouping the number of RAI papers by topic, we observed that companies were also leading in certain areas. In Accountability, Amazon and Google were among the top institutions, along with Stanford University and Oxford University; in Explainability, Microsoft shares the top places with TU Berlin; and in Fairness, Google leads along with NUS and CMU.
@@
RAI papers going into patents or repositories have unconventional combinations of prior work. In general, papers going into patents or repositories are grounded on unconventional combinations of prior work. However, Privacy papers that went into patents can be more conventional (Figure 5a), similar to Sustainability papers that went into repositories (Figure 5b). To see this distinction in a concrete example, consider two unconventional papers: “Alleviating Privacy Attacks via Causal Learning” by Tople, Sharma, and Nori (2020) and “Practical One-Shot Federated Learning for Cross-Silo Setting” by Li, He, and Song (2021). These papers draw from uncommon pairs of venues such as NeurIPS and IEEE Computer Security Foundations Symposium, and Conference on Computer and Communications Security and ICLR and have consequently gone into several repositories. While security and machine learning papers might be more similar in their approach, let us consider another example of two unusual papers in Accountability: “Data-Centric Explanations: Explaining Training Data of Machine Learning Systems to Promote Transparency” by Anik and Bunt (2021) and “Amortized Generation of Sequential Algorithmic Recourses for Black-Box Models” by (Verma, Hines, and Dickerson 2022). These papers draw from uncommon pairs of venues such as HCI and ICML. On the other hand, we found that Sustainability papers going into repositories are more likely to be grounded on conventional combinations of prior work.
@@ Discussion
By analyzing 1,747 papers in the field of RAI from leading Computer Science conferences, we found that a small subset that goes into patents or repositories is highly cited. The majority of these papers that go into patents focus on sustainability and privacy, while those that go into repositories focus on explainability. The temporal dynamics between papers that go into patents and those that go into repositories vary. The time frame for papers that go into patents is 8 years, while those that go into repositories are 1 year. Our survival analysis revealed not only the presence of a rapid publication cycle and a slower patent production cycle, but also that these cycles vary across different RAI topics. This suggests the need to reevaluate existing assumptions about the pace and consistency of knowledge production in both patents and repositories. When examining the top institutions producing impactful RAI papers in terms of patents and repositories, we observed that the impact of top U.S. institutions is not exclusive. Although our study concentrated on patents in the U.S., we noticed a significant contribution from European and Asian institutions. This indicates a more global distribution of influence than previously recognized. Finally, RAI papers generally rely on unconventional knowledge, often integrating expertise from fields such as ML and HCI, in both patents and repositories.
@@
By synthesizing our findings with those documented in prior literature, we propose three recommendations for enhancing the translational landscape of RAI research.
@@
Developing Initiatives Between Academia and Industry. Translational research is crucial in bridging the gap between scientific breakthroughs and their practical application (Searles et al. 2016). Echoing the findings of previous studies (Cao et al. 2023), our study suggests that only a portion of RAI research significantly contributes to patents and repositories. Therefore, it is beneficial to foster collaborative efforts between academia and industry. This can be achieved through various initiatives. First, by establishing dual mentoring programs that connect RAI researchers with mentors from both academic and industrial sectors, including patent professionals. Second, by organizing joint conferences and patent clinics that bring together AI researchers, patent engineers, and industry experts, researchers can gain immediate access to professional guidance on patenting. Finally, by promoting corporate participation in academic endeavors through hackathons and sponsorships (Briscoe and Mulligan 2014), with the goal of enhancing the translation of RAI research into practical innovations.
@@
Balancing and Diversifying Focus in RAI Research. The observed delay between RAI papers and patents may suggest an overemphasis on certain research topics. While this focus on specific areas (e.g., Fairness and Accountability) advances theoretical knowledge, it might inadvertently slow down the development of practical innovations (McKinsey & Company 2022). However, slowing down is not necessarily bad. In fact, incremental approaches could be favorable when considering the impact of technical work on social change (Abebe et al. 2020). This is advised because otherwise researchers may fail to recognize how changes in behaviors of pre-existing social systems can be caused by the introduction of technology. It can also lead to technical solutions focused purely on ethical or value-aligned deployments, without considering whether a given system functions and provides benefits (Raji et al. 2022), or even a technical solution is needed at all (Selbst et al. 2019). To address these challenges, a more strategic approach is necessary to accelerate the transition from academic discoveries to patentable technologies. This would involve reevaluating RAI research priorities to ensure they align with practical needs. Our findings also suggest that papers that go into patents tend to concentrate on sustainability and privacy, whereas those in repositories are focused on explainability. This discrepancy highlights the importance for academia to realign its research priorities with industry demands.
@@
Fostering Cross-discipline RAI Research. The impact of RAI papers is shaped by how they build upon prior knowledge. Generally, RAI papers tend to use unconventional combinations of prior knowledge, and often draw upon expertise from diverse fields such as ML and HCI. To continue making a significant impact on both innovation and development, it is important to encourage researchers to engage in cross-disciplinary research efforts. This can be achieved by the formation of research teams that include experts from multiple disciplines (Rakova et al. 2021) (e.g., computer science, social sciences, and law), incorporating diverse perspectives and technical expertise (Moitra et al. 2022) as well as contributing towards organizational culture shift (Rakova et al. 2021). Additionally, educational programs that promote interdisciplinary work (Frodeman 2011) are required to train AI researchers in ethical, legal, and social implications of their work (McDonald and Pan 2020).
@@ Conclusion
By analyzing 1,747 RAI papers, we found that RAI research, despite being an emerging field, contributes to patents and development. This translation process typically occurs within a time frame of 1 to 8 years, highlighting a gap between research output and translational outcomes. Our findings also challenge the conventional notion that impactful RAI research is predominantly the domain of top U.S. institutions. Instead, European and Asian institutions are also major contributors, indicating a more globally distributed expertise in the field. Finally, the multidisciplinary nature of RAI papers, which often build upon unconventional combinations of prior knowledge, is vital for ensuring the development of fair, transparent, and accountable AI systems.
Researcher Positionality Statement and Ethical Considerations
The research team includes three men from Southeast Asia and Southern Europe, working in industry. Our shared backgrounds include Human-Computer Interaction, software engineering, AI, social and ubiquitous computing, and urbanism. In light of recent calls for RAI research to include impact and ethical statements (Olteanu et al. 2023), our research neither involves any subjects nor do we foresee any significant harm that could result.
@@ Appendix A Keywords to Filter RAI Papers
Removed keywords. These keywords were removed from our query because they were non-discriminative as they are often used to describe generally AI. The 19 keywords included: validation, reliability, correctness, accuracy, robustness, generalizability, authenticity, quality, measurability, dependability, capability, safety, security, resilience, confidentiality, integrity, availability, usability, controllability.
@@ Appendix B Analyzed Venues
Figure A.1: The cumulative number of citations in millions as a function of top venues considered in each research area. The top 10 venues in the six research areas (AI, Database, NLP, HCI, Data Mining, and Computer Vision) cover more than 88% of the total citations in each area. The numbers at the top of each subplot show the total coverage of citations from the top 10 venues in each area.
@@
Table A.1: Distribution of papers (including non-RAI papers) over the years for each research area. Venues are ranked based on their h5 index impact factor. According to Google Scholar, h5 index is the h-index for articles published in the last five complete years; it is the largest number $h$ such that $h$ articles published in 2018-2022 have at least h citations each.